Given the vertex of a parabola of (4, –8) that passes through the point (2, 20), what is the vertex form of the quadratic equation?
Found 2 solutions by josgarithmetic, TeachMath:
Answer by josgarithmetic(39617)

y=a(x-4)^2-8, if this parabola has vertical symmetry axis
or
x=a(x-4)+8, if parabola has horizontal symmetry axis

Solve the equation for a, and substitute the x & y coordinates of the point, and evaluate your factor a.

Answer by TeachMath(96)

Vertex form: y = a(x - h)^2 + k
20 = a(2 - 4)^2 + - 8
20 = 4a - 8__4a = 28___a = 28/4 = 7

Equation: y = 7(x - 4)^2 - 8